OpenStudy (anonymous):

Two x-interecepts of the function h(x) = -2sin(x- π/2) are

OpenStudy (experimentx):

solve for x -2sin(x- π/2) = 0

OpenStudy (anonymous):

with the sin in there do i do sin-1 instead of regular sin to figure it out?

OpenStudy (campbell_st):

the curve have been shifted right by a distance of pi/2 the normal root for -2sin(x) is the origin, then pi, 2pi etc the phase shift right will see the intercepts move to -pi/2, pi2, 3pi/2, 5pi/2 etc

OpenStudy (experimentx):

or, \[ \sin ( x - \pi/2) = 0 \] or, \[ \sin(x- \pi/2) = \sin 0 = \sin \pi = \sin 2\pi = \sin 3pi = \sin 4\pi = ... \] \[ x-\pi/2 = 0 = \pi = 2\pi = ... \]

OpenStudy (anonymous):

π/3 and 5π/2 would that be the answer

OpenStudy (experimentx):

find any two values of x

OpenStudy (anonymous):

nvm answer would be 0 and 2pi?

OpenStudy (experimentx):

plugin those values... into woflramalpha, and see if you get 0 or not